Write the sum as a product: sin(15.12) - sin(10.92) = Preview
Answer #1

\sin(15.1x)-\sin(10.9x)=2\cos(\frac{15.1x+10.9x}{2})\sin(\frac{15.1x-10.9x}{2}) \\ \\ =2\cos(\frac{26x}{2})\sin(\frac{4.2x}{2}) \\ \\ =2\cos(13x)\sin(2.1x)

[sum as a product

\sin(A)-\sin(B)=2\cos(\frac{A+B}{2})\sin(\frac{A-B}{2})]